# Nearly geodesic surfaces are filling

### v2

@@ 3.3 Nearly geodesic surfaces are strongly filling
[AA23] and the proofs of Theorem 1.6 and Theorem 1.1 will reveal the subtle difference between measure convergence and Hausdorff convergence. It is illuminating to examine hyperbolic surfaces first. Let $S$ be a genus-$2$ hyperbolic surface and $\beta$ a simple close geodesic. By $D_{\beta}$ we mean applying the Dehn twist with respect to $\beta$. Let $\mathcal{PML}$ denote the space of projective measured laminations on $S$ and $C(S)$ denote the collection of closed sets in $S$ equipped with Hausdorff metric. From the basic theory on geodesic laminations (e.g., [Mar16, 8.3]), we have as $n\rightarrow\infty$,
$\frac{1}{2n}D_{\beta_{1}}^{n}D_{\beta_{2}}^{n}\gamma$ $\rightarrow\frac{1}{2}\beta_{1}\sqcup\frac{1}{2}\beta_{2}\textnormal{ in }\mathcal{PML}\textnormal{ with the weak-}^{*}\textnormal{ topology,}$
@@
In contrast, every geodesic plane in a closed hyperbolic $3$-manifold $M$ is either dense or a closed Fuchsian surface by Ratner’s and Shah’s theorem [Rat91, Sha91a]. In Al Assal’s thesis [AA23], he proves that the set of weak- limits of the probability area measures induced on the Grassmann bundle by asymptotically Fuchsian minimal surfaces achieves every convex combination of ergodic $SL_{2}(\mathbb{R})$-invariant measures. When $M$ has no Fuchsian surfaces, the limiting measure must be the Liouville measure which implies that $\mathcal{G}_{2}S_{i}$ equidistribute. When $M$ has a Fuchsian surface, Theorem 1.6 suggests that there is a dense subset in the Hausdorff limit with measure 0. In particular, to connect distinct (or identical) components in the Hausdorff limit, we must connect by a dense subset, unless the asymptotically Fuchsian surfaces belong to the same commensurability class (in which case the geodesic planes in $\mathbb{H}^{3}$ or $\mathcal{G}_{2}M$ which cover them coincide).
@@
The existence of a smooth minimal surface representative follows from Schoen-Yau [SY79, Lemma 3.2 and Theorem 2.1] and Sacks-Uhlenbeck [SU82]. Osserman [Oss70] and Gulliver [Gul73] prove such minimal surfaces have no branched points. The smoothness of the tangent bundle of $S$ implies that $\mathcal{G}_{2}S$ is a connected closed subset of $\mathcal{G}_{2}M$. The uniqueness of the minimal surface representative in 2.3 allows us to identify a homotopy class $[S]$ of $(1+\epsilon)$-quasi-Fuchsian surfaces with a unique subset $\mathcal{G}_{2}S$ of $\mathcal{G}_{2}M$. ∎
[AA23] proves that for a sequence of asymptotically Fuchsian surfaces $S_{i}$, both the pleated maps $f^{p}_{i}$ and minimal maps $f^{m}_{i}$ homotopic to $S_{i}\hookrightarrow M$ have the same weak- measure limit along any convergent subsequence. The universal cover of $f^{p}_{i}$ may develop wrinkles and each $f^{p}_{i}$ lifts to a disconnected subset of $\mathcal{G}_{2}M$.
@@
Let $M$ be a closed hyperbolic $3$-manifold. If $S_{i}$ is a sequence of asymptotically Fuchsian surfaces in distinct commensurability classes, then all but finitely many $S_{i}]$ are strongly filling.
@@
Let $C(\mathcal{G}_{2}M)$ denote the set of closed subsets in $\mathcal{G}_{2}M$, equipped with Hausdorff metric. Since $M$ is closed, $\mathcal{G}_{2}M$ is closed, and hence $C(\mathcal{G}_{2}M)$ is a compact metric space. Thus $\mathcal{G}_{2}S_{i}$ has a convergent subsequence, which we still denote by $\mathcal{G}_{2}S_{i}$. By 3.10, each $\mathcal{G}_{2}S_{i}$ is a connected closed subset of $\mathcal{G}_{2}M$, and thus
@@
Here $S$ is either one or uncountably many geodesic planes in $\mathbb{H}^{3}$.
@@
If the limit $\mathcal{G}_{2}S$ is not a single geodesic plane which is discrete, it is dense and $\mathcal{G}_{2}S_{i}\rightarrow\mathcal{G}_{2}M$.
@@
Enumerate the collection of discrete geodesic planes as $\{P_{i}\}$ which are embedded and mutually disjoint in $\mathcal{G}_{2}M$. Denote
$\mathcal{P}_{j}$ $\coloneqq\{P_{i}:i\leq j\},$
$\mathcal{P}_{L}$ $\coloneqq\{\textnormal{geodesic planes which are dense in }\mathcal{G}_{2}M\}.$
For any finite $j$, the union $\cup_{\mathcal{P}_{j}}P_{i}$ is a disconnected closed set in $\mathcal{G}_{2}M$. Hence $\mathcal{G}_{2}S\cap\cup_{\mathcal{P}_{j}}P_{i}$ is a strict subset of $\mathcal{G}_{2}S$ for any $j$, unless $\mathcal{G}_{2}S\subset P_{i}$ for some $i$, contradiction. Thus for any genus $g$, there exist planes $P_{k}$ of genus >g>g such that $P_{k}\subset\mathcal{G}_{2}S-\cup_{\mathcal{P}_{j}}P_{i}$ by 3.12. Since $P_{k}$ equidistribute as $k\rightarrow\infty$ by [MS95], $P_{k}\subset\mathcal{G}_{2}S$ pointed converge to a dense plane $P\in\mathcal{P}_{L}$. Since $\mathcal{G}_{2}S$ is closed and $PSL_{2}(\mathbb{R})$-invariant, $P\subset\mathcal{G}_{2}S$. ∎
Now suppose that $\mathcal{G}_{2}S$ is supported on a single discrete geodesic plane $P$. Then for any small $\eta>0$, $\mathcal{G}_{2}S_{i}\subset N_{\eta}(\mathcal{G}_{2}S)$ for all sufficiently large $i$.
@@
Let $S\subset M$ be totally geodesic and $\eta$ smaller than the normal injectivity radius of $\mathcal{G}_{2}S$ in $\mathcal{G}_{2}M$. If $S^{\prime}$ is smoothly immersed $\pi_{1}$-injective and $\mathcal{G}_{2}S^{\prime}\subset N_{\eta}(\mathcal{G}_{2}S)$, then $S^{\prime}$ and $S$ have homotopic images.
@@
By a homotopy, $S^{\prime}$ is minimal. Since $\mathcal{G}_{2}S$ and $N_{\eta}(\mathcal{G}_{2}S)$ are embedded, the preimages of $N_{\eta}(\mathcal{G}_{2}S)$ in $\mathcal{G}_{2}\mathbb{H}^{3}$ consist of pairwise disjoint embedded neighborhoods $\{N_{\eta}(P)|\textnormal{Pr}(P)\subset\mathbb{H}^{3}\textnormal{ cover }S\}$ (where $P$ is a totally geodesic plane). By 3.10, $\mathcal{G}_{2}S^{\prime}$ is connected, and hence lifts to $\mathcal{G}_{2}\widetilde{S}^{\prime}\subset N_{\eta}(P)$. The projection $\textnormal{Pr}:\mathcal{G}_{2}\mathbb{H}^{3}\rightarrow\mathbb{H}^{3}$ maps $\mathcal{G}_{2}\widetilde{S}^{\prime}$ to $\widetilde{S}^{\prime}\subset N_{\eta}(\textnormal{Pr}(P))$. The only complete minimal disk contained in $N_{\eta}(\textnormal{Pr}(P))$ is the totally geodesic plane $\textnormal{Pr}(P)$ itself. Thus $\widetilde{S}^{\prime}=\textnormal{Pr}(P)$ and $S^{\prime}$ is commensurable with $S$. ∎
@@
If a sequence of Grassmann bundles $\mathcal{G}_{2}S_{i}$ over minimal surfaces converges to a dense subset of $\mathcal{G}_{2}M$, then all but finitely many $S_{i}$ are filling.
@@
We just proved there is a subsequence $S_{j}$ of the initial sequence $S_{i}$ which is strongly filling. If there exist infinitely many non-filling $S_{i}$, then we apply the above argument again to this collection and conclude it must contain a subsequence consisting of filling surfaces, which is a contradiction. ∎
3.13 establishes Theorem 1.6. Lowe in [Low21, Theorem 4.2] proves that the asymptotically Fuchsian surfaces are dense in $\mathcal{G}_{2}M$ assuming they have a limiting circle which is not the limit set of a closed Fuchsian surface. Fernando Al Assal and Ben Lowe [AL25] independently obtain Theorem 1.6 for finite-volume hyperbolic manifolds (including the noncompact case) and obtain various applications, including rigidity results for negatively curved manifolds [AL25, Theorem 1.7] and a gap theorem [AL25, Theorem 1.6] for the quasi-Fuchsian constants of nearly geodesic surfaces in geometrically finite hyperbolic manifolds of infinite volume.
@@
Al Assal in [AA23] constructs asymptotically Fuchsian surfaces which equidistribute or whose limit has a nontrivial scarring component. Building upon Liu-Markovic [LM15], in [AA23, Section 4, 5, and 7], in order to construct connected asymptotically Fuchsian surfaces with various limiting properties, Al Assal uses all $(R,\epsilon)$-good pants which are known to equidistribute (different limiting measures correspond to different weights on the set of good pants). Thus 3.13 and [AL25] imply as $\epsilon$ gets smaller, the pants required to construct a connected $(1+\epsilon)$-quasi-Fuchsian surface become dense.
@@
Since the proof of Theorem 1.6 relies on the fact that asymptotically Fuchsian surfaces eventually become dense in $\mathcal{G}_{2}M$, this naturally implies that they form a ubiquitous set.
The proof of 1.9.
@@
The idea is essentially in the proof of 3.15. Using Theorem 1.6, the Grassmann bundles $\mathcal{G}_{2}S_{i}$ converge to $\mathcal{G}_{2}M$. Let $C_{1},C_{2}$ be two disjoint and non tangent circles at $\partial\widetilde{M}$ which bound two geodesic planes $P_{1},P_{2}\subset\mathbb{H}^{3}$, respectively. We want to show there exists some $S_{i}$ which has a preimage $\tilde{S}_{i}$ whose limit set separates $C_{1}$ and $C_{2}$.
@@
Then the second paragraph of the proof of 3.15 implies that for sufficiently large $i$, $\partial\tilde{S}_{i}$ separates $C_{1}$ and $C_{2}$ and $\tilde{S}_{i}$ separate $P_{1}$ and $P_{2}$. ∎
@@ 3.4 The dichotomy of embedded surfaces in a hyperbolic $3$-manifold
The incompressibility of $S$ and $S^{\prime}$ induces an injection of the fundamental group:
@@
Suppose $S$ is embedded in $M$. Then, by lifting to the covering $M^{\prime}$ of $M$ corresponding to $\pi_{1}(S^{\prime})$, we also lift $S$ homeomorphically to an embedded incompressible surface $\tilde{S}$ in $M^{\prime}$. By the classification of $I$-bundles over surfaces [Hem76, 10.6 Theorem], we have $M^{\prime}\cong S^{\prime}\times I$. By the classification of incompressible embedded surfaces in an $I$-bundle [Mar16, Proposition 9.3.18], $\tilde{S}$ is isotopic to the horizontal surface $S^{\prime}\times 0$. This contradicts the assumption that $\tilde{S}$ is a nontrivial covering of $S^{\prime}$. ∎
Proof of 1.2.
@@
Let $S$ be an embedded surface in $M$. By 3.17, $S$ is a primitive orientable surface. By the Lefschetz duality (e.g., [Mar16, Corollary 9.1.5]), we have $\dim H^{1}(M-S;\mathbb{Q})>0$ which implies that $S$ is never filling. By Theorem 1.1, there exists an $\epsilon_{0}>0$ such that every $(1+\epsilon)$-quasi-Fuchsian surface with $0<\epsilon\leq\epsilon_{0}$ is filling. Thus an embedded non Fuchsian surface satisfies $\epsilon>\epsilon_{0}$.
Now let $S\hookrightarrow M$ be an embedded minimal surface of supremum curvature $\lambda_{0}\leq\frac{\epsilon_{0}}{2+\epsilon_{0}}<1$. We want to show $S$ is totally geodesic of genus $g_{0}$. By Uhlenbeck [Uhl83], having supremum curvature <1<1 implies that $S$ is almost-Fuchsian, say $(1+\epsilon)$-quasi-Fuchsian. By Theorem 2.2,
$1+\epsilon<\frac{1+\lambda_{0}}{1-\lambda_{0}}\leq 1+\epsilon_{0}.$
If $\epsilon\neq 0$, then Theorem 1.1 forces $S$ to be filling and hence non embedded. Thus $\epsilon=0$ and $S$ is Fuchsian. Suppose the genus $\mathbf{g}(S)\geq g_{0}$. If $S$ is an orientable surface attaining the smallest genus in its commensurability class, then Theorem 1.8 implies $S$ is filling. If $S$ is commensurable with a Fuchsian surface of genus $g_{0}$, since commensurable Fuchsian surfaces span the same subset of $M$ (i.e., geometrically identical or same $PSL_{2}(\mathbb{R})$-orbit), the immersion $S\hookrightarrow M$ maps some points in $S$ to points in $M$ with multiplicity more than $1$. This violates the assumption that $S$ is embedded. Thus $S$ must have genus $g_{0}$. ∎
[Miy94, Theorem 4.2] bounds the number of embedded Fuchsian surfaces in $M$ by the volume of $M$ up to a uniform constant. Let $\mathcal{E}_{0}$ be the set of embedded Fuchsian surfaces. There are many non-Haken manifolds $M$ in which case $\mathcal{S}^{E}=\emptyset$. A uniform quasi-Fuchsian constant lower bound suggests most embedded surfaces are some definite amount away from totally geodesic. This can also be measured geometrically using a uniform lower bound on the supremum principle curvature $\lambda_{0}$, i.e.,
@@
when we take $S$ to be minimal in its homotopy class. The principal curvature lower bound uses [HW13, Theorem 1.2]. A fibered surface is uniformly away from being totally geodesic by considering either the supremum principal curvature [Uhl83] or the Hausdorff dimension of its limit set.

### v3

@@ 3.4 Nearly geodesic surfaces are strongly filling
[AA23] and the proofs of Theorem 1.4 and Theorem 1.1 will reveal the subtle difference between measure convergence and Hausdorff convergence. It is illuminating to examine hyperbolic surfaces first. Let $S$ be a genus-$2$ hyperbolic surface and $\beta$ a simple closed geodesic. By $D_{\beta}$ we mean applying the Dehn twist with respect to $\beta$. Let $\mathcal{PML}$ denote the space of projective measured laminations on $S$ and $C(S)$ denote the collection of closed sets in $S$ equipped with Hausdorff metric. From the basic theory of geodesic laminations (e.g., [Mar16, 8.3]), we have as $n\rightarrow\infty$,
$\frac{1}{\ell(D_{\beta_{1}}^{n}D_{\beta_{2}}^{n}\gamma)}D_{\beta_{1}}^{n}D_{\beta_{2}}^{n}\gamma$ $\rightarrow\frac{1}{2}\beta_{1}\sqcup\frac{1}{2}\beta_{2}\textnormal{ in }\mathcal{PML}\textnormal{ with the weak-}^{*}\textnormal{ topology,}$
@@
In contrast, every plane in a closed hyperbolic $3$-manifold $M$ is either dense or a closed Fuchsian surface by Ratner’s and Shah’s theorem [Rat91, Sha91a]. In Al Assal’s thesis [AA23], he proves that the set of weak- limits of the probability area measures induced on the Grassmann bundle by asymptotically geodesic minimal surfaces achieves every convex combination of ergodic $PSL_{2}(\mathbb{R})$-invariant measures. When $M$ has no Fuchsian surfaces, the limiting measure must be the Liouville measure which implies that $\mathcal{G}_{2}S_{i}$ equidistribute. When $M$ has a Fuchsian surface $S$, Theorem 1.4 says that even if the limiting measure of a sequence of asymptotically geodesic surfaces $S_{i}$ is entirely concentrated on $S$, $\mathcal{G}_{2}S_{i}$ is still asymptotically dense in $\mathcal{G}_{2}M$, unless $S_{i}$ are eventually commensurable with $S$. If the limiting measure of $S_{i}$ is not entirely supported on a single Fuchsian surface, then $\mathcal{G}_{2}S_{i}$ must be asymptotically dense.
@@
The existence of a smooth minimal surface representative follows from Schoen-Yau [SY79, Lemma 3.2 and Theorem 2.1] and Sacks-Uhlenbeck [SU82]. Osserman [Oss70] and Gulliver [Gul73] prove such minimal surfaces have no branched points. The smoothness of the tangent bundle of $S$ implies that $\mathcal{G}_{2}S$ is a connected closed subset of $\mathcal{G}_{2}M$. The uniqueness of the minimal surface representative in Theorem 2.3 allows us to identify a homotopy class $[S]$ of $(1+\epsilon)$-quasi-Fuchsian surfaces with a unique subset $\mathcal{G}_{2}S$ of $\mathcal{G}_{2}M$. ∎
[AA23] proves that for a sequence of asymptotically geodesic surfaces $S_{i}$, both the pleated maps $f^{p}_{i}$ and minimal maps $f^{m}_{i}$ homotopic to $S_{i}\hookrightarrow M$ have the same weak- measure limit along any convergent subsequence. The universal cover of $f^{p}_{i}$ may develop wrinkles and each $f^{p}_{i}$ lifts to a disconnected subset of $\mathcal{G}_{2}M$.
@@
Let $M$ be a closed hyperbolic $3$-manifold. If $S_{i}$ is a sequence of classes of asymptotically geodesic surfaces, then all but finitely many $S_{i}]$ are strongly filling.
@@
Let $C(\mathcal{G}_{2}M)$ denote the set of closed subsets in $\mathcal{G}_{2}M$, equipped with Hausdorff metric. Since $M$ is closed, $\mathcal{G}_{2}M$ is closed, and hence $C(\mathcal{G}_{2}M)$ is a compact metric space. Thus $\mathcal{G}_{2}S_{i}$ has a convergent subsequence, which we still denote by $\mathcal{G}_{2}S_{i}$. By 3.12, each $\mathcal{G}_{2}S_{i}$ is a connected closed subset of $\mathcal{G}_{2}M$, and thus
@@
Here $S$ is either one or uncountably many planes in $\mathbb{H}^{3}$.
@@
If the limit $\mathcal{G}_{2}S$ is not a single discrete plane, it is dense and $\mathcal{G}_{2}S_{i}\rightarrow\mathcal{G}_{2}M$.
@@
Enumerate the collection of discrete planes as $\{P_{i}\}$ which are embedded and mutually disjoint in $\mathcal{G}_{2}M$. Denote $\mathcal{P}_{j}\coloneqq\{P_{i}:i\leq j\}.$ By 3.14, $\mathcal{G}_{2}S$ is $PSL_{2}(\mathbb{R})$-invariant, and thus $\mathcal{G}_{2}S\cap P_{i}$ is either $P_{i}$ or the empty set. For any finite $j$, the union $\cup_{\mathcal{P}_{j}}P_{i}$ is a disconnected closed set in $\mathcal{G}_{2}M$. Hence $\mathcal{G}_{2}S\cap\cup_{\mathcal{P}_{j}}P_{i}$ is a strict subset of $\mathcal{G}_{2}S$ for any $j$, unless $\mathcal{G}_{2}S\subset P_{i}$ for some $i$, contradiction. Thus for any genus $g$, there exist planes $P_{k}$ of genus >g>g such that $P_{k}\subset\mathcal{G}_{2}S-\cup_{\mathcal{P}_{j}}P_{i}$ by 3.14. Since $P_{k}$ equidistribute as $k\rightarrow\infty$ by [MS95], $P_{k}\subset\mathcal{G}_{2}S$ Hausdorff converge to a plane $P$ dense in $\mathcal{G}_{2}M$. Since $\mathcal{G}_{2}S$ is closed and $PSL_{2}(\mathbb{R})$-invariant, $P\subset\mathcal{G}_{2}S$. ∎
Now suppose that $\mathcal{G}_{2}S$ is supported on a single discrete plane $P$. Then for any small $\eta>0$, $\mathcal{G}_{2}S_{i}\subset N_{\eta}(\mathcal{G}_{2}S)$ for all sufficiently large $i$.
@@
Let $S\subset M$ be totally geodesic and $\eta$ smaller than the normal injectivity radius of $\mathcal{G}_{2}S$ in $\mathcal{G}_{2}M$. If $S^{\prime}$ is a smoothly immersed $\pi_{1}$-injective surface and $\mathcal{G}_{2}S^{\prime}\subset N_{\eta}(\mathcal{G}_{2}S)$, then $S^{\prime}$ and $S$ have homotopic images.
@@
By a homotopy, $S^{\prime}$ is minimal. Since $\mathcal{G}_{2}S$ and $N_{\eta}(\mathcal{G}_{2}S)$ are embedded, the preimages of $N_{\eta}(\mathcal{G}_{2}S)$ in $\mathcal{G}_{2}\mathbb{H}^{3}$ consist of pairwise disjoint embedded neighborhoods $\{N_{\eta}(P)|\textnormal{Pr}(P)\subset\mathbb{H}^{3}\textnormal{ cover }S\}$ (where $P\subset\mathcal{G}_{2}\mathbb{H}^{3}$ is a plane). By 3.12, $\mathcal{G}_{2}S^{\prime}$ is connected, and hence lifts to $\mathcal{G}_{2}\widetilde{S}^{\prime}\subset N_{\eta}(P)$. The projection $\textnormal{Pr}:\mathcal{G}_{2}\mathbb{H}^{3}\rightarrow\mathbb{H}^{3}$ maps $\mathcal{G}_{2}\widetilde{S}^{\prime}$ to $\widetilde{S}^{\prime}\subset N_{\eta}(\textnormal{Pr}(P))$. The only complete minimal disk contained in $N_{\eta}(\textnormal{Pr}(P))$ is the plane $\textnormal{Pr}(P)$ itself. Thus $\widetilde{S}^{\prime}=\textnormal{Pr}(P)$ and $S^{\prime}$ is commensurable with $S$. ∎
@@
If a sequence of Grassmann bundles $\mathcal{G}_{2}S_{i}$ over minimal surfaces converges to a dense subset of $\mathcal{G}_{2}M$, then all but finitely many $S_{i}$ are strongly filling.
@@
As in the second paragraph of the proof of 3.14, the limit of the pointed Hausdorff convergence of $(\widetilde{p_{j}},\tilde{S}_{j})$ is the same as graphical convergence on compact sets to a totally geodesic disc, after passing to a subsequence.
@@
We just proved there is a subsequence $S_{j}$ of the initial sequence $S_{i}$ which is strongly filling. If there exist infinitely many non-filling $S_{i}$, then we apply the above argument again to this collection and conclude it must contain a subsequence consisting of strongly filling surfaces, which is a contradiction. ∎
[…]
@@
[…]
@@
The proof of Theorem 1.4 implies that the boundary circles of asymptotically geodesic surfaces in $\partial\mathbb{H}^{3}$ forms a ubiquitous set.
The proof of 1.9 (ubiquity of the boundary circles of asymptotically geodesic surfaces).
@@
The idea is essentially in the proof of 3.17. Using Theorem 1.4, the Grassmann bundles $\mathcal{G}_{2}S_{i}$ converge to $\mathcal{G}_{2}M$. Let $C_{1},C_{2}$ be two disjoint and non tangent circles at $\partial\widetilde{M}$ which bound two planes $P_{1},P_{2}\subset\mathbb{H}^{3}$, respectively. We want to show there exists some $S_{i}$ which has a preimage $\tilde{S}_{i}$ whose limit set separates $C_{1}$ and $C_{2}$.
@@
Then the second paragraph of the proof of 3.17 implies that for sufficiently large $i$, $\partial\tilde{S}_{i}$ separates $C_{1}$ and $C_{2}$ and $\tilde{S}_{i}$ separate $P_{1}$ and $P_{2}$. ∎
@@ 3.5 The dichotomy of embedded surfaces in a hyperbolic $3$-manifold
The $\pi_{1}$-injectivity of $S$ and $S^{\prime}$ in $M$ induces an injection of the fundamental group:
@@
Suppose $S$ is embedded in $M$. Then, by lifting to the covering $M^{\prime}$ of $M$ corresponding to $\pi_{1}(S^{\prime})$, we also lift $S$ homeomorphically to an embedded $\pi_{1}$-injective surface $\tilde{S}$ in $M^{\prime}$. Since $\pi_{1}(M^{\prime})\cong\pi_{1}(S^{\prime})$, by the classification of $3$-manifolds whose fundamental groups are virtually surface subgroups ([Hem76, 10.6 Theorem]), we have $M^{\prime}\cong S^{\prime}\times I$. By the classification of $\pi_{1}$-injective embedded surfaces in an $I$-bundle [Mar16, Proposition 9.3.18], $\tilde{S}$ is isotopic to the horizontal surface $S^{\prime}\times 0$. This contradicts the assumption that $\tilde{S}$ is a nontrivial covering of $S^{\prime}$. ∎
Proof of 1.2 (on the dichotomy of embedded $\pi_{1}$-injective surfaces in $M$).
@@
Let $S$ be an embedded surface in $M$. By 3.19, $S$ is a primitive orientable surface. By the Lefschetz duality (e.g., [Mar16, Corollary 9.1.5]), we have $\dim H^{1}(M-S;\mathbb{Q})>0$ which implies that $S$ is never filling. By Theorem 1.1, there exists an $\epsilon_{0}>0$ such that every $(1+\epsilon)$-quasi-Fuchsian surface with $0<\epsilon\leq\epsilon_{0}$ is filling. Thus an embedded non Fuchsian surface satisfies $\epsilon>\epsilon_{0}$.
Now let $S\hookrightarrow M$ be an embedded minimal surface of supremum curvature $\lambda_{0}\leq\frac{\epsilon_{0}}{2+\epsilon_{0}}<1$. We want to show $S$ is totally geodesic of genus $\mathfrak{g}_{0}$. By Uhlenbeck [Uhl83], having supremum curvature <1<1 implies that $S$ is almost-Fuchsian, say $(1+\epsilon)$-quasi-Fuchsian. By Theorem 2.2,
$1+\epsilon\leq\frac{1+\lambda_{0}}{1-\lambda_{0}}\leq 1+\epsilon_{0}.$
If $\epsilon\neq 0$, then Theorem 1.1 forces $S$ to be filling and hence non embedded. Thus $\epsilon=0$ and $S$ is Fuchsian. Suppose the genus $\mathfrak{g}[S]\geq\mathfrak{g}_{0}$. If $S$ is an orientable surface attaining the smallest genus in its commensurability class, then Theorem 1.1 implies $S$ is filling. If $S$ is commensurable with a Fuchsian surface of genus $\mathfrak{g}_{0}$, since commensurable Fuchsian surfaces span the same subset of $M$ (i.e., geometrically identical or same $PSL_{2}(\mathbb{R})$-orbit), the immersion $S\hookrightarrow M$ maps some points in $S$ to points in $M$ with multiplicity more than $1$. This violates the assumption that $S$ is embedded. Thus $S$ must have genus $\mathfrak{g}_{0}$. ∎
The following example from Dunfield gives infinitely many closed hyperbolic $3$-manifolds with infinitely many embedded quasi-Fuchsian surfaces.
Example 3.20.
Let $K$ be a Montesinos knot with at least four rational tangles with all $q_{i}\geq 3$. By [Oer84, Corollary 3], $M=S^{3}-K$ contains embedded $\pi_{1}$-injective surfaces of every genus >1>1. Let $Y_{n}$ be $1/n$ Dehn surgery on $M$, and pick $n$ sufficiently large so that $Y_{n}$ is closed hyperbolic by Thurston’s Dehn surgery theorem (e.g., [Mar16, Corollary 15.1.3]). By the proof of [Oer84, Corollary 4(b)], all the surfaces in $M$ remain essential in $Y_{n}$. Since $H_{1}(Y_{n};\mathbb{Z})$ = 0, none of these surfaces can be fibers or semi-fibers. Bonahon-Thurston’s theorem [Thu78, Bon86] says that $\pi_{1}$-injective surfaces in $M$ is either quasi-Fuchsian or virtually fibered. Hence, $Y_{n}$ is a closed hyperbolic $3$-manifold with essential embedded quasi-Fuchsian surfaces of every genus >1>1.
[Miy94, Theorem 4.2] bounds the number of embedded Fuchsian surfaces in $M$ by the volume of $M$ up to a uniform constant. Let $\mathcal{E}_{0}$ be the set of embedded Fuchsian surfaces and $\mathcal{E}_{\geq\epsilon_{0}}$ the set of quasi-Fuchsian surfaces with constant at least $1+\epsilon_{0}$ but not any lower. The deviation from being totally geodesic can also be measured geometrically using a uniform lower bound on the supremum principal curvature $\lambda_{0}$, i.e.,
@@
when we take $S$ to be minimal in its homotopy class. The principal curvature lower bound uses [Eps86, (5.5)]. A fibered surface is uniformly away from being totally geodesic by considering either the supremum principal curvature [Uhl83] or the Hausdorff dimension of its limit set.